OpenStudy (anonymous):

if g equals 3 what is the value of 2g plus fifty-three

OpenStudy (anonymous):

2g + 53 = ? (subst. 3 in there)

OpenStudy (anonymous):

59

OpenStudy (anonymous):

2*3= 6 6+53=59

OpenStudy (anonymous):

Good job! :)